NERSCPowering Scientific Discovery Since 1974

April 27

PDSF Users Meeting 4/27/10

Attending: Eric and Jay from PDSF and users Marjorie, Jeff P, Andrei and Craig.

Cluster status: Utilization has been light.

Outages: chos problem on interactives has been fixed (java vm can use a lot of memory).

Upcoming downtimes: in June/July will have multiple days center-wide outage - concerns were expressed about the timing - Jay to send out an email.

New hardware: Waiting for Dell to start selling new Intel chips. ATLAS storage is up and running. 40 new ALICE nodes in rack 21 and 22.

ATLAS grid activities. A validation test required some bestman patches.

Other Topics:

- STAR testing on /project: need to define next set of tests.

- Jeff noted that the eliza1io limit is still zero

- Dayabay using SL4 but trying SL5. X11 headers missing?

- STAR won't need 32-bit SL5 chos but is testing 64-bit SL5 build

- /common inodes got used up yesterday, STAR and ATLAS cleaning up.

- In September DOE will come to visit ATLAS. They want to spend the money they have before then.